# A Novel Algorithm for Digital Lithological Mapping: Case Studies in Sri Lanka’s Mineral Exploration

### v1

@@ 3.3 Ilmenite
The proposed algorithm is used to identify the presence of Ilmenite in two geographical locations in Sri Lanka. Pulmoddai in the northeastern coastal region of the island is considered to be one of the richest mineral sand depositsEkanayake et al., (2019). This sand reserve is known for Ilmenite, Rutile, Leucoxene, Garnet, Monazite, Zircon, and sillimanite as well as minor amounts of Magnetite (Amalan et al., (2018)). The soil samples were collected from the sites near the coastline as well as in the vicinity of the distributors of Yan Oya. As the second region of interest, the southern coast of Mannar Island which is located on the northwest coastline of Sri Lanka was selected. Mannar has been identified as a potential area for heavy mineral placer deposits. It was observed that the geographical regions constituted mainly of vegetation, soil, paddy fields, beaches with the sand, sea, and other water bodies.
@@ 3.4.1 Hyperion Sensor and Its Details
The HSIs required for the study were captured through the Hyperion Sensor of “NASA’s Earth Observing 1” satellite. The spectral resolution of the instrument is 242 bands ranging from 0.4 to 2.5 um with a spatial resolution of 30 meters. For the data acquisition, the United States Geological Survey (USGS) database was used. The acquired images are represented in Fig. 1. Conducting the analysis on the original data strip requires high computational power. Therefore, the regions of interest were cropped from the original HSI to minimize the computational burden.
Figure 1: Study Locations and Satellite Paths over the locations. The regions illustrated are : (a) Mannar (b) Murunkan (c) Jaffna (d) Pulmoddai
All the images captured from the Hyperion sensor are provided as Level 1Gst products which are terrain corrected and available in the form of 16-bit radiance values. In order to perform further analysis on the product the radiance values were converted into reflectance values according to the equation proposed by (Kokaly et al., (2017)). Ancillary data required for this purpose such as mean solar exo-atmospheric irradiance for each band, and earth-sun distance in astronomical units for a set of days of the year are available in the USGS database. The geometrical parameters used for the conversion are given in Table 2.
@@ 3.4.2 Extraction of Laboratory Reference Signature for each Mineral
The proposed algorithm requires a laboratory spectral signature of the explored mineral to use as the reference. These spectral signatures for the minerals; Ilmenite, Montmorillonite, and Limestone were obtained from the USGS Spectral Library (Kokaly et al., (2017)). However, the laboratory spectral signatures of the minerals have been obtained through different sensing equipment. Therefore, an interpolation approach was carried out to promote continuity among the spectral bands. The details of the three instruments are tabulated in Table 3.
@@
Sensor Wavelength (um) Number of Channels Mineral
@@
Analytic Spectral Devices Full Range 0.35 - 2.5 2151 Limestone
@@ 4.1 Pre Classification
Remotely sensed HSIs contain both pure pixels that correspond to specific types of materials (such as water, vegetation, soil, sand, etc.) as well as mixed pixels. The specific endmembers of an HSI are influenced by the topography of the selected environment (Guo et al., (2018)) while the choice of the number of endmembers depends on the application and the complexity of hyperspectral data. Most of the algorithms used in the literature for endmember extraction, require prior knowledge of the number of endmembers in the hyperspectral data (Zare and Gholizadeh, (2019)). The methods of identifying the number of endmembers present in a scene include visual inspection (Ekanayake et al., (2019)) and statistical analysis (Sobrino and Jiménez-Muñoz, (2014)).
@@
Once the number of dominant endmembers has been identified, the HSI is pre-classified into sub-components. For this, Vertex Component Analysis (VCA) is used. First, the candidate endmembers were extracted using VCA for the number of classes suggested by the Elbow method. Then, each HSI pixel was normalized in the spectral direction using the L2-norm. Similarly, the candidate endmember signatures were also normalized. Thereafter, the Euclidean distance between the HSI spectra and each candidate end member was calculated. The final goal of this is to isolate the pixels corresponding to each candidate end member. Therefore, the reciprocal of the computed distance metric was used as a measure of affinity. The formula used for the similarity measure is computed through the following equation.
@@ 4.2 Subclass Identification and Representative Endmember Extraction
Once the pixels of the HSI are classified into the respective endmember classes based on their spectral properties, it can be seen that the classification perfectly highlights the land cover mapping of the given terrain (Fig. 2). For the mineral representative pixel identification, the algorithm prefers the extraction of a mineral representative signature endemic to the scene from the HSI itself. As the targeted minerals are mostly present in soil, initially, the soil pixels were isolated from the others. The isolation of soil pixels before mineral-based classification reduces errors that arise due to misclassification, confusion due to mixed pixels and reduces the computational cost. Next, the algorithm tries to identify the highly correlated pixels to the actual spectral signature of the target mineral. For this purpose, the laboratory mineral signatures for the three minerals Limestone, Montmorillonite and, Ilmenite were acquired from the USGS spectral library (Kokaly et al., (2017)). The sensor details of the spectral library signatures and the mandatory preprocessing steps taken are mentioned in 3.4.2.
@@
Figure 2: Generated Abundance Maps after pre-classification: (a) Jaffna (b) Pulmoddai (c) Mannar (d) Giant’s Tank
@@
where, $r_{k}$, $x_{k,n}$, $s_{n}$, and $b$ represent the Pearson correlation coefficient between $k^{th}$ pixel and the reference signature, value of the $n^{th}$ spectral band of the $k^{th}$ pixel, value of the $n^{th}$ spectral band of the laboratory Limestone signature, and the number of spectral bands, respectively.
In order to identify the highly correlated and the slightly correlated pixel classes with the mineral, threshold values for classification should be computed. Even though the computed correlation coefficients show the relationship between the soil signature with the ideal mineral signature, it might not reflect the relationship between the signature endemic to the specific site. However, this can be taken as an initial estimation for identifying pixels which tend to show closeness to the mineral at the site. Based on this, it was decided to create two classes, namely, the mineral representative class and the impurity representative class based on the computed correlation coefficients. In order to do this, it is necessary to identify the threshold values used to classify the pixels into the aforementioned classes.To compute these threshold values, initially, two endmembers were extracted from the soil pixel manifold and the correlation of these pixels to the laboratory reference was computed. This approach was based on the hypothesis that, depending on the availability of the mineral in the area covered by a pixel, the pixel’s signature would exhibit inherent features. The authors try to extract two endmembers from the soil pixels using VCA. As VCA tries to separate two source signatures from the soil pixels, it was assumed that the mineral’s presence could be a differentiating factor when it comes to the extraction of the end members.
The signature which showed a high correlation to the reference mineral signature was chosen to be the mineral representative signature while the signature with a low correlation was assigned to represent the impurities naming it as the impurity representative signature. The threshold values for classification were taken to be the correlation coefficients of the above representative pixels.
  • 1.
@@
The pixels that showed a higher correlation than $C_{1}$ were assigned to the mineral representative sub-class, while the pixels that showed a lower correlation than $C_{2}$ were assigned to the impurity representative sub-class. Finally, the class representative signatures which were extracted from VCA prior to assigning the pixels into subclasses were replaced by the mean signature of each sub-class. The steps taken are summarized through the pseudo code.
@@
8:Create ${soil\_matrix}$ by selection soil pixels from the HSI
9:Extract 2 endmembers performing VCA on ${soil\_mat}$ and store as columns of ${rep\_sig}$ matrix.
@@
20:     if $corr(x)>upper$ then
21:         Label $HSI(pixel)$ as ${}^{\prime}mineral_{r}epresentative^{\prime}$
22:     else if $corr(x)<lower$ then
23:         Label $HSI(pixel)$ as ${}^{\prime}impurityrepresentative^{\prime}$
@@
25:         Label $HSI(pixel)$ as ${}^{\prime}notasoilpixel^{\prime}$
@@ 4.3 Alignment of soil pixels for purity
To identify the mineral availability of the particular terrain, and map the relative abundances of the earmarked mineral, it is necessary to investigate the correlation between the soil pixel and the mineral reference signature. This will allow one to identify patterns in the variation of correlation when traversing from the least correlated set to the most correlated set along the soil signature manifold. To quantify the comparative availability of the mineral in the particular region of interest, we define a parameter named Relative Availability (RA)(Equation 3). This parameter is a normalized metric to measure the amount of mineral in a given soil pixel. In order to compute Relative Availability, the authors utilize a dimensionality reduction method. In the context of dimensionality reduction, several algorithms such as Principal component analysis, Singular Value Decomposition, t-SNE, and Fisher’s Discriminant Analysis (FDA) have been widely used(Pearson, (1901); Golub and Reinsch, (1970); Xu et al., (2004)).
However, to analyze the changing patterns of correlation when moving along the soil signature manifold conditioned on the mineral availability, it is possible to project the dataset onto a direction where it enhances the separation between the aforementioned classes. FDA is a statistical technique used for dimensionality reduction while promoting discrimination among the classes. It tries to identify an eigen-direction that maximizes the separation between classes. To be concise, it aims to reduce the projected within-class variation and increase the gap between projected means at the same time. Within this framework, FDA was used to identify an eigen-direction in which there is a clear separation between the two sub-classes.
Initially, FDA was performed on the two representative classes acquired in the previous step. Through this, the eigen-direction with the highest separation between the two classes was identified. Then, the pixel vectors belonging to the soil subclass were projected onto the new direction. Further, representative signatures of the two classes were also transformed onto the reduced one-dimensional space following the calculation of RA through the equation given below. Then, the distances to both representative signatures from the pixel were calculated for each pixel along the eigen-direction. The reciprocal of each distance was taken as a measurement (Kizel et al., (2017)) for the affinity and the relative availability of the mineral was defined as,
@@ 4.4 Least Square Estimation for Abundance Generation
After the calculation of relative abundance values for the mineral of interest, the authors use least squares estimation to further enhance the calculated mineral abundance assuming a linear mixture model between the mineral representative signature and the impurity representative signature. Initially, the mineral representative signature was replaced by the average signature of the pixels emphasizing a relative availability greater than 0.8. Similarly, the impurity signature was replaced by the average signature of the pixels with a relative availability lower than 0.2. Assuming a linear mixture model the soil pixel signature, $\mathbf{s}$ can be represented as,
@@
where $\mathbf{m}$ and $\mathbf{\alpha}$ denote the mineral representative signature and its coefficient, while $\mathbf{r}$ and $\mathbf{\beta}$ represent the impurity representative signature and its coefficient. The coefficients represent the fractions of each of the components in the soil pixel signature. By utilizing this formulation, the non-negative least squares estimation was used to calculate the optimal coefficients, $\mathbf{\alpha^{*}}$ and $\mathbf{\beta^{*}}$, of the linear mixture model.
In order to formulate the estimation of coefficients as a non-negative least squares problem, each pixel $\mathbf{s}$ is considered separately and estimated using the $\mathbf{A}$ matrix which contains the two representative signatures (as column vectors). Furthermore, the vector $\mathbf{x}$ contains the two coefficients, $\mathbf{\alpha}$ and $\mathbf{\beta}$ as its elements. Therefore, the optimal vector $\mathbf{x^{*}}$ is found by minimizing the objective function with the constraints given below.
@@ 4.5 Field Survey and Sample Collection
As outlined in section 3, samples of Ilmenite, Limestone, and Montmorillonite were gathered from four distinct regions within Sri Lanka. After the algorithm had generated likely mineral availability locations, the authors, in coordination with the Department of Geology at the University of Peradeniya, Sri Lanka, conducted the site selection process. This selection was meticulously carried out, taking into consideration regions with both high and low mineral availability, with the aim of validating the algorithm’s performance.
The mineral map generated in this study served as a valuable tool for identifying potential areas with both high and low probabilities of containing the desired minerals. Once the pixels for the survey were selected from the generated mineral map, the precise GPS coordinates for the corresponding pixels were determined using ArcGIS. As it is impractical to collect samples from such a large area, the sample collection was carried out with the assumption that the soil is of a homogeneous composition. The representations of the selected sites for the sample collection are depicted in Figures 8 (c), 9 (c), 11 (c), 12 (c).
Upon arriving at the sites of interest, soil samples were meticulously collected under the supervision of experienced members from the Department of Geology at the University of Peradeniya, Sri Lanka. Samples were obtained from various locations scattered across the sites of interest. During the collection process, extra care was taken to gently scrape the surface in order to gather samples without digging into the soil. This precaution is essential because remote sensing (RS) images solely capture material or information available on the surface.
@@ 4.6 Magnetic Separation and XRD for Mineral Identification
To examine the presence of the earmarked minerals the samples obtained were put under laboratory testing. For the identification of the minerals X-ray diffraction (XRD) test and Magnetic separation was carried out. As appear in the literature, the preferred testing procedure for Limestone and Montmorillonite (Wanasinghe and Adikary, (2017)) is XRD while for Ilmenite it is magnetic separation. However, due to the texture and grain size of the samples collected from Mannar, it was decided to use the XRD test instead the magnetic separation only for finding the availability of Ilmenite in these soil samples. The mineral identification for the samples collected from Pulmoddai was done through magnetic separation using the FRANTZ Magnetic Separator Model L-1. Before testing for the mineral availability each sample was prepared adhering to the standard procedure with the aid of the Department of Geology, Faculty of Science, University of Peradeniya. The procedure followed is described in the following sections.

### v2

@@ 3.3 Ilmenite
The proposed algorithm is used to identify the presence of Ilmenite in two geographical locations in Sri Lanka. Pulmoddai in the northeastern coastal region of the island is considered to be one of the richest mineral sand deposits(Ekanayake et al. 2019). This sand reserve is known for Ilmenite, Rutile, Leucoxene, Garnet, Monazite, Zircon, and Sillimanite, as well as minor amounts of Magnetite (Amalan et al. 2018). The soil samples were collected from the sites near the coastline as well as in the vicinity of the distributors of Yan Oya. As the second region of interest, the southern coast of Mannar Island which is located on the northwest coastline of Sri Lanka was selected. Mannar has been identified as a potential area for heavy mineral placer deposits. It was observed that the geographical regions constituted mainly of vegetation, soil, paddy fields, beaches with the sand, sea, and other water bodies.
@@ 3.4.1 Hyperion Sensor and Its Details
The HSIs required for the study were captured through the Hyperion Sensor of “NASA’s Earth Observing 1” satellite. The spectral resolution of the instrument is 242 bands ranging from 0.4 to 2.5 um with a spatial resolution of 30 meters. For the data acquisition, the United States Geological Survey (USGS) database was used. The acquired images are represented in Fig. 1. Conducting the analysis on the original data strip requires high computational power. Therefore, the regions of interest were cropped from the original HSI to minimise the computational burden.
Figure 1: Study Locations
All the images captured from the Hyperion sensor are provided as Level 1Gst products which are terrain-corrected and available in the form of 16-bit radiance values. In order to perform further analysis on the product the radiance values were converted into reflectance values according to the equation proposed by (Kokaly et al. 2017). Ancillary data required for this purpose such as mean solar exo-atmospheric irradiance for each band, and earth-sun distance in astronomical units for a set of days of the year are available in the USGS database. The geometrical parameters used for the conversion are given in Table 2.
@@ 3.4.2 Extraction of Laboratory Reference Signature for each Mineral
The proposed algorithm requires a laboratory spectral signature of the explored mineral to use as the reference. These spectral signatures for the minerals; Ilmenite, Montmorillonite, and Limestone were obtained from the USGS Spectral Library (Kokaly et al. 2017). However, the laboratory spectral signatures of the minerals have been obtained through different sensing equipment. Therefore, an interpolation approach was carried out to promote continuity among the spectral bands. The details of the three instruments are tabulated in Table 3.
@@
Sensor Wavelength ($\mu$m) Number of Channels Mineral
@@
Analytic Spectral 0.35 - 2.5 2151 Limestone
Devices Full Range
@@ 4.1 Pre Classification
Remotely sensed HSIs contain both pure pixels that correspond to specific types of materials (such as water, vegetation, soil, sand, etc.) as well as mixed pixels. The specific endmembers of an HSI are influenced by the topography of the selected environment (Guo et al. 2018) while the choice of the number of endmembers depends on the application and the complexity of hyperspectral data. Most of the algorithms used in the literature for endmember extraction, require prior knowledge of the number of endmembers in the hyperspectral data (Zare and Gholizadeh 2019). The methods of identifying the number of endmembers present in a scene include visual inspection (Ekanayake et al. 2019) and statistical analysis (Sobrino and Jiménez-Muñoz 2014).
@@
Once the number of dominant endmembers has been identified, the HSI is pre-classified into sub-components. For this, Vertex Component Analysis (VCA) is used. First, the candidate endmembers were extracted using VCA for the number of classes suggested by the Elbow method. Then, each HSI pixel was normalized in the spectral direction using the L2-norm. Similarly, the candidate endmember signatures were also normalised. Thereafter, the Euclidean distance between the HSI spectra and each candidate end member was calculated. The final goal of this is to isolate the pixels corresponding to each candidate end member. Therefore, the reciprocal of the computed distance metric was used as a measure of affinity. The formula used for the similarity measure is computed through the following equation.
@@ 4.2 Subclass Identification and Representative Endmember Extraction
Once the pixels of the HSI are classified into the respective endmember classes based on their spectral properties, it can be seen that the classification perfectly highlights the land cover mapping of the given terrain (Fig. 2). For the mineral representative pixel identification, the algorithm prefers the extraction of a mineral representative signature endemic to the scene from the HSI itself. As the targeted minerals are mostly present in soil, initially, the soil pixels were isolated from the others. The isolation of soil pixels before mineral-based classification reduces errors that arise due to misclassification, and confusion due to mixed pixels and reduces the computational cost. Next, the algorithm tries to identify the highly correlated pixels to the actual spectral signature of the target mineral. For this purpose, the laboratory mineral signatures for the three minerals Limestone, Montmorillonite, and Ilmenite were acquired from the USGS spectral library (Kokaly et al. 2017). The sensor details of the spectral library signatures and the mandatory preprocessing steps taken are mentioned in 3.4.2.
@@
(d) Giant’s Tank
Figure 2: Generated Abundance Maps after pre-classification
@@
where, $r_{k}$, $x_{k,n}$, $s_{n}$, and $b$ represent the Pearson correlation coefficient between $k^{th}$ pixel and the reference signature, the value of the $n^{th}$ spectral band of the $k^{th}$ pixel, the value of the $n^{th}$ spectral band of the laboratory Limestone signature, and the number of spectral bands, respectively.
In order to identify the highly correlated and the slightly correlated pixel classes with the mineral, threshold values for classification should be computed. Even though the computed correlation coefficients show the relationship between the soil signature with the ideal mineral signature, it might not reflect the relationship between the signature endemic to the specific site. However, this can be taken as an initial estimation for identifying pixels that tend to show closeness to the mineral at the site. Based on this, it was decided to create two classes, namely, the mineral representative class and the impurity representative class based on the computed correlation coefficients. In order to do this, it is necessary to identify the threshold values used to classify the pixels into the aforementioned classes. To compute these threshold values, initially, two endmembers were extracted from the soil pixel manifold, and the correlation of these pixels to the laboratory reference was computed. This approach was based on the hypothesis that, depending on the availability of the mineral in the area covered by a pixel, the pixel’s signature would exhibit inherent features. The authors try to extract two endmembers from the soil pixels using VCA. As VCA tries to separate two source signatures from the soil pixels, it was assumed that the mineral’s presence could be a differentiating factor when it comes to the extraction of the end members.
The signature that showed a high correlation to the reference mineral signature was chosen to be the mineral representative signature while the signature with a low correlation was assigned to represent the impurities naming it as the impurity representative signature. The threshold values for classification were taken to be the correlation coefficients of the above representative pixels.
@@
The pixels that showed a higher correlation than $C_{1}$ were assigned to the mineral representative sub-class, while the pixels that showed a lower correlation than $C_{2}$ were assigned to the impurity representative sub-class. Finally, the class representative signatures which were extracted from VCA prior to assigning the pixels into subclasses were replaced by the mean signature of each sub-class. The steps taken are summarized through the pseudo-code.
@@
8:Create ${soil\_matrix}$ by selecting soil pixels from the HSI
9:Extract 2 endmembers by performing VCA on ${soil\_mat}$ and store as columns of ${rep\_sig}$ matrix.
@@
20:     if $corr(pixel)>upper$ then
21:         Label $HSI(pixel)$ as $mineral\_representative$
22:     else if $corr(pixel)<lower$ then
23:         Label $HSI(pixel)$ as $impurity\_representative$
@@
25:         Label $HSI(pixel)$ as $not\_a\_soil\_pixel$
@@ 4.3 Alignment of soil pixels for purity
To identify the mineral availability of the particular terrain and map the relative abundances of the earmarked mineral, it is necessary to investigate the correlation between the soil pixel and the mineral reference signature. This will allow one to identify patterns in the correlation variation when traversing from the least correlated set to the most correlated one along the soil signature manifold. To quantify the comparative availability of the mineral in the particular region of interest, we define a parameter named Relative Availability (RA)(Equation 3). This parameter is a normalised metric to measure the amount of minerals in a given soil pixel. In order to compute Relative Availability, the authors utilise a dimensionality reduction method. In the context of dimensionality reduction, several algorithms such as Principal component analysis, Singular Value Decomposition, t-SNE, and Fisher’s Discriminant Analysis (FDA) have been widely used(Pearson 1901; Golub and Reinsch 1970; Xu, Yang, and Jin 2004).
However, to analyse the changing patterns of correlation when moving along the soil signature manifold conditioned on the mineral availability, it is possible to project the dataset in a direction that enhances the separation between the aforementioned classes. FDA is a statistical technique used to reduce dimensionality while promoting class discrimination. It tries to identify an eigen-direction that maximises the separation between classes. To be concise, it aims to reduce the projected within-class variation and simultaneously increase the gap between projected means. Within this framework, FDA was used to identify an eigen-direction with a clear separation between the two sub-classes.
Initially, FDA was performed on the two representative classes acquired in the previous step. Through this, the eigen-direction with the highest separation between the two classes was identified. Then, the pixel vectors belonging to the soil subclass were projected onto the new direction. Further, representative signatures of the two classes were also transformed onto the reduced one-dimensional space following the calculation of RA through the equation given below. Then, the distances to both representative signatures from the pixel were calculated for each pixel along the eigen-direction. The reciprocal of each distance was taken as a measurement (Kizel et al. 2017) for the affinity, and the relative availability of the mineral was defined as,
@@ 4.4 Least Square Estimation for Abundance Generation
After calculating relative abundance values for the mineral of interest, the authors use least squares estimation to further enhance the calculated mineral abundance, assuming a linear mixture model between the mineral representative signature and the impurity representative signature. Initially, the mineral representative signature was replaced by the average signature of the pixels, emphasizing a relative availability greater than 0.8. Similarly, the impurity signature was replaced by the average signature of the pixels with a relative availability lower than 0.2. Assuming a linear mixture model the soil pixel signature, $\mathbf{s}$ can be represented as,
@@
where $\mathbf{m}$ and $\mathbf{\alpha}$ denote the mineral representative signature and its coefficient, while $\mathbf{r}$ and $\mathbf{\beta}$ represent the impurity representative signature and its coefficient. The coefficients represent the fractions of each component in the soil pixel signature. By utilizing this formulation, the non-negative least squares estimation was used to calculate the optimal coefficients, $\mathbf{\alpha^{*}}$ and $\mathbf{\beta^{*}}$, of the linear mixture model.
In order to formulate the estimation of coefficients as a non-negative least squares problem, each pixel $\mathbf{s}$ is considered separately and estimated using the $\mathbf{A}$ matrix, which contains the two representative signatures (as column vectors). Furthermore, the vector $\mathbf{x}$ contains the two coefficients, $\mathbf{\alpha}$ and $\mathbf{\beta}$ as its elements. Therefore, the optimal vector $\mathbf{x^{*}}$ is found by minimizing the objective function with the constraints given below.
@@ 4.5 Field Survey and Sample Collection
As outlined in section 3, samples of Ilmenite, Limestone, and Montmorillonite were gathered from four distinct regions within Sri Lanka. After the algorithm had generated likely mineral availability locations, the authors, in coordination with the Department of Geology at the University of Peradeniya, Sri Lanka, conducted the site selection process. This selection was meticulously carried out, considering regions with both high and low mineral availability, aiming to validate the algorithm’s performance.
The mineral map generated in this study served as a valuable tool for identifying potential areas with high and low probabilities of containing the desired minerals. Once the pixels for the survey were selected from the generated mineral map, the precise GPS coordinates for the corresponding pixels were determined using ArcGIS. As it is impractical to collect samples from such a large area, the sample collection was carried out with the assumption that the soil is of a homogeneous composition. The representations of the selected sites for the sample collection are depicted in Figures 8 (c), 9 (c), 11 (c), 12 (c).
Upon arriving at the sites of interest, soil samples were meticulously collected under the supervision of experienced members from the Department of Geology at the University of Peradeniya, Sri Lanka. Samples were obtained from various locations scattered across the sites of interest. During the collection process, extra care was taken to gently scrape the surface to gather samples without digging into the soil. This precaution is essential because remote sensing (RS) images solely capture material or information available on the surface.
@@ 4.6 Magnetic Separation and XRD for Mineral Identification
The samples obtained were put under laboratory testing to examine the presence of the earmarked minerals. X-ray diffraction (XRD) test and Magnetic separation were carried out to identify the minerals. As the literature shows, the preferred testing procedure for Limestone and Montmorillonite (Wanasinghe and Adikary 2017) is XRD, while for Ilmenite, it is magnetic separation. However, due to the texture and grain size of the samples collected from Mannar, it was decided to use the XRD test instead of the magnetic separation only to find the availability of Ilmenite in these soil samples. The mineral identification for the samples collected from Pulmoddai was done through magnetic separation using the FRANTZ Magnetic Separator Model L-1. Before testing for the mineral availability, each sample was prepared to adhere to the standard procedure with the aid of the Department of Geology, Faculty of Science, University of Peradeniya. The procedure followed is described in the following sections.